OpenStudy (anonymous):

PLEASE HELP! MEDAL WILL BE REWARDED! FIND THE DERIVATIVE: f(x)=secx-sqrt2 tanx Is the answer sec^2(x)[sin(x) - √2] or f(x) = sec(x) - (sqrt(2) * tanx)???

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what's the derivative of sec(x)?

OpenStudy (anonymous):

its secx tanx

jimthompson5910 (jim_thompson5910):

what is the derivative of tan(x)?

OpenStudy (anonymous):

sec^2x

jimthompson5910 (jim_thompson5910):

and just so I have it correct, the original function is \[\large f(x) = \sec(x) - \sqrt{2}*\tan(x)\] right? and the tan(x) is NOT in the square root right?

OpenStudy (anonymous):

yes you are correct

jimthompson5910 (jim_thompson5910):

so that would mean that we would get this \[\large f(x) = \sec(x) - \sqrt{2}*\tan(x)\] \[\large \frac{d}{dx}[f(x)] = \frac{d}{dx}[\sec(x) - \sqrt{2}*\tan(x)]\] \[\large \frac{d}{dx}[f(x)] = \frac{d}{dx}[\sec(x)] - \frac{d}{dx}[\sqrt{2}*\tan(x)]\] \[\large f^{\prime}(x) = \sec(x)\tan(x) - \sqrt{2}*\sec^2(x)\] \[\large f^{\prime}(x) = \sec(x)\left[\tan(x) - \sqrt{2}*\sec(x)\right]\] Note: the last step is entirely optional really (since it's just factoring out the GCF)

OpenStudy (anonymous):

Thank you so much!!!!

jimthompson5910 (jim_thompson5910):

sure thing